OpenStudy (anonymous):

Solve the equation: log(base4)x+log(base4)(x-3)=1

OpenStudy (anonymous):

\[log_4(x)+log_4(x-3)=1\] rewrite as a single logarithm using property of the logs that \[log(ab)=log(a)+log(b)\] but use it from right to left \[log_4(x(x-3))=1\] rewrite in equivalent exponential form using \[log_b(x)=y\] means \[b^y=x\] \[x(x-3)=4^1\] solve the quadratic equation \[x^2-3x-4=0\] \[(x-4)(x-1)=0\] \[x=4\] or \[x=-1\] but you cannot take log of a negative number so only 4 is the answer

OpenStudy (anonymous):

thanks!